OpenStudy (anonymous):

Express the radical in simplest form: (squ 98x)

OpenStudy (jdoe0001):

can you simplify 98?

OpenStudy (anonymous):

yes it ends in a even number

OpenStudy (jdoe0001):

... what factors did you get for it?

OpenStudy (anonymous):

2*49=98 7*14=98

OpenStudy (jdoe0001):

 so we can say that \(\bf 98 = 2 \times 7 \times 7 \implies 2 \times 7^2\\\quad \\ \sqrt{98x} \implies \sqrt{2 \times 7^2x } \implies 7\sqrt{2x}\)